All questions go with Model 4 , please label thank you
in the Ltwis suettre for PCls and XeFe Model 4: Which Lewis Structure is Better? The two structures below could represent the thiocyanate polyatomic ion. However only one does. Formal charges can be used to distinguish which has a lower energy and is therefore a better description of the lowest energy state of this ion. How is formal charge determined? 9. 10. How is formal charge used to identify unreasonable Lewis structures? What are the formal charges on the atoms in the thiocyanate ion in model 4? Write the formal charge in parentheses above each atom in these structures. 11.
12. Which structure in Model 4 has the lower energy and is therefore the better description of the lowest energy state of this ion? Model 5: NO2- The Need for Multiple Lewis Structures (Resonance) tructures which have the same distribution of low formal
Answer #1

9) Formal charge = Valence electrons-lone pair electron-bonds formed

10) Higher the formal charge ,lesser the significance of structure. A structure in which positive charge is present on electropositive atom and negative charge on electronegative atom is more significant.

11) S-C\equivN-

Formal charge on S = 6-6-1 = -1

Formal charge on C = 4-0-4 = 0

Formal charge on N = 5-2-3 = 0

C\equivN-S-

Formal charge on C = 4-2-3 = -1

Formal charge on N = 5-0-4 = 1

Formla charge on S = 6-6-1 = -1

12) Structure one has low energy because it has least formal charge.